# A ‘Catch, Tag, and Release’ Mechanism for Embeddings

### v1

@@ 4.1 Setup and Main Result
where the $\operatorname{softmax}(\cdot)$ is causal.
{mdframed}
[shadow=true, shadowsize=1pt, innertopmargin=10pt, innerbottommargin=10pt]
Theorem 4.1 (‘Catch, Tag, and Release’ Theory).
Assume the learnable parameters, $\theta$, satisfy:
$s_{\texttt{num}}=0$
${\bm{W}}^{1}_{V}=\begin{bmatrix}0&0\\ 0&-1\end{bmatrix}$
${\bm{W}}^{2}_{Q}{\bm{W}}^{2}_{K}=\begin{bmatrix}0&b\\ 0&d\end{bmatrix},\ \ \text{where}\ \ d>0,b\in\mathbb{R}$
${\bm{W}}^{2}_{V}=\begin{bmatrix}1\\ 0\end{bmatrix}$ (3)
Then:
$f_{\theta}({\bm{x}})\stackrel{{\scriptstyle s_{\texttt{tag}}\to\infty}}{{\xrightarrow{\hskip 28.45274pt}}}\frac{1}{T-t}\sum_{i=t+1}^{T}x_{i}$
for any $x_{1},...,x_{T}\in\mathbb{R}$. The model will have the following features for any sequence ${\bm{x}}$:
  • the  [SEP] forms an attention sink,
  • the tokens after the  [SEP] have activation outliers in their second coordinate,
  • the second coordinate is an outlier feature dimension, and
  • the attention matrices ${\bm{W}}^{1}_{V}$, ${\bm{W}}^{2}_{Q}$, ${\bm{W}}^{2}_{K}$, ${\bm{W}}^{2}_{V}$ are low rank, projecting either to the tag or number subspace
Remark 4.2.
The model utilizes a tagging mechanism based on the second coordinate of the token representations. A token is classified as:
  Untagged:
if its second coordinate is finite.
  Tagged:
if its second coordinate diverges.
Initially, all number tokens are assumed to be untagged, with their tag coordinate set to $-1$.
Figure 6: Attention probabilities in the first attention layer. The [SEP] token acts as an attention sink for tokens ${\bm{x}}_{i}$ with $i\geq t$. Consistent with Equation (4), the remaining probabilities in the same rows decay exponentially as $\mathcal{O}(e^{-s_{\texttt{tag}}})$. Notably, the attention sink does not influence the probabilities of tokens ${\bm{x}}_{i}$ for $i<t$, which remain $\mathcal{O}(1)$.
4.2 Proof of Theorem, Part 1: ‘Catch, Tag, Release’
@@
Thus, [SEP]acts as an attention sink, where all tokens $x_{i}$ for $i>t$, as well as the [SEP] token itself, attend exclusively to the [SEP]token. As a result, the attention of all tokens $x_{i}$ for $i>t$ and the [SEP] token have been caught. These observations are depicted in Figure 6 above.
@@
[…]
@@
4.3 Proof of Theorem, Part 2: Leveraging the Tags
@@
[…]
@@
[…]
@@
Combining all three cases together with Equation (7) leads to:
$\lim_{s_{\texttt{tag}}\to\infty}{\bm{A}}^{2}_{T,j}=\begin{cases}0,&j<t\\ 0,&j=t\\ \frac{1}{T-t},&j>t\\ \end{cases}$
The above shows how the tags have been leveraged to arrive at a uniform distribution over the desired tokens. Then using this, Equation (6), and Equation (3), we can conclude that:
@@
4.4 Proof of Theorem, Part 3: Proving the Three Cases
@@
where the order of magnitude of the last two terms is given by Equation (4) and depicted in Figure 6.
@@
which we show in Appendix A.
@@
where $\lim_{s_{\texttt{tag}}\to\infty}({\bm{A}}^{1}_{j,t}-{\bm{A}}^{1}_{T,T})s_{\texttt{tag}}=0$ is proved in Appendix A.
5 Related Work
Register Buffers in Transformers
Darcet et al. (2024) observed that vision transformers exhibit high-norm outlier tokens, similar to LLMs, and that registers (data-independent tokens) are needed to prevent them from arising. This aligns closely with the perspective proposed by Sun et al. (2024a); Gu et al. (2024) that the first token attention sink is serving as a bias term for the attention mechanism.
Adam Causes Sinks and Outliers
Another recent avenue of investigation is whether attention sinks and outlier feature dimensions are a by-product of the optimizer. Both Kaul et al. (2024) and Guo et al. (2024b), show that the Adam optimizer (Kingma and Ba, 2015) leads to attention sinks and outlier features. In the former, they propose OrthoAdam which applies a rotation to the gradients to prevent any specific outlier dimensions.
Rank Collapse
Recent studies have shown that stacking self-attention layers in transformers can lead to rank collapse, where token representations lose dimensionality due to inherent architectural properties (Noci et al., 2022; Geshkovski et al., 2024). This phenomenon may stem from the ‘catch, tag, release’ mechanism, in which attention concentrates around attention sinks, causing tokens to cluster around them and become confined to a low-rank subspace.
Low-Rank Terms for Model Compression
Beyond OATS, other works have also proposed leveraging low-rank terms to mitigate compression loss for deep networks (Yu et al., 2017; Mozaffari and Dehnavi, 2024; Li et al., 2025). Another line of research is to incorporate a low-rank adapter during compression that is fine-tuned to mitigate drops in performance (Li et al., 2023; Zhang et al., 2023; Guo et al., 2024a; Zhao et al., 2024; Mozaffari et al., 2025).
6 Conclusion
This paper establishes that low-rank structures in attention weight matrices cause attention sinks and outlier features, which, in turn, enable the ‘catch, tag, and release’ mechanism. This mechanism serves as a fundamental organizational principle in transformers, dynamically segmenting and tagging token sequences to facilitate tasks such as few-shot learning and chain-of-thought reasoning. Empirically, we demonstrate that pruning techniques that fail to preserve low-rank structures disrupt this mechanism, significantly impairing few-shot generalization. Through rigorous theoretical analysis, we prove that ‘catch, tag, and release’ is not incidental but necessary for performing fundamental tasks like averaging a sub-sequence of numbers. This provides a compelling explanation for why this phenomenon emerges organically in transformer architectures.
Acknowledgements
This research was supported in part by the Province of Ontario, the Government of Canada through CIFAR, and industry sponsors of the Vector Institute (www.vectorinstitute.ai/partnerships/current-partners/). This research was also enabled in part by support provided by Compute Ontario (https://www.computeontario.ca) and the Digital Research Alliance of Canada (https://alliancecan.ca).
Appendix A Proof of Auxiliary Limits
@@
$\lim_{s_{\texttt{tag}}\to\infty}s_{\texttt{tag}}^{2}({\bm{A}}^{1}_{t,t}-{\bm{A}}^{1}_{T,t})=\lim_{s_{\texttt{tag}}\to\infty}\frac{\sum_{k=1,k\neq t}^{T}s_{\texttt{tag}}^{2}\exp(x_{T}x_{k}+1-s_{\texttt{tag}})-(t-1)s_{\texttt{tag}}^{2}\exp(s_{\texttt{tag}}-s_{\texttt{tag}}^{2})}{\left(1+(t-1)\exp(s_{\texttt{tag}}-s_{\texttt{tag}}^{2})\right)\cdot\left(1+\sum_{k=1,k\neq t}^{T}\exp(x_{T}x_{k}+1-s_{\texttt{tag}})\right)}=\frac{0}{1}=0$
@@
$\lim_{s_{\texttt{tag}}\to\infty}({\bm{A}}^{1}_{j,t}-{\bm{A}}^{1}_{T,T})s_{\texttt{tag}}$ $=\frac{\sum_{k=1,k\neq t}^{T}s_{\texttt{tag}}\exp(x_{T}x_{k}+1-s_{\texttt{tag}})-\sum_{k=1,k\neq t}^{j}s_{\texttt{tag}}\exp(x_{j}x_{k}+1-s_{\texttt{tag}})}{\left(1+\sum_{k=1,k\neq t}^{j}\exp(x_{j}x_{k}+1-s_{\texttt{tag}})\right)\cdot\left(1+\sum_{k=1,k\neq t}^{T}\exp(x_{T}x_{k}+1-s_{\texttt{tag}})\right)}$
@@
$\lim_{s_{\texttt{tag}}\to\infty}({\bm{A}}^{1}_{j,t}-{\bm{A}}^{1}_{T,T})s_{\texttt{tag}}^{2}$ $=\frac{\sum_{k=1,k\neq t}^{T}s_{\texttt{tag}}^{2}\exp(x_{T}x_{k}+1-s_{\texttt{tag}})-\sum_{k=1,k\neq t}^{j}s_{\texttt{tag}}^{2}\exp(x_{j}x_{k}+1-s_{\texttt{tag}})}{\left(1+\sum_{k=1,k\neq t}^{j}\exp(x_{j}x_{k}+1-s_{\texttt{tag}})\right)\cdot\left(1+\sum_{k=1,k\neq t}^{T}\exp(x_{T}x_{k}+1-s_{\texttt{tag}})\right)}$
@@
Appendix B Further Experimental Results
(a) Input embedding visualization at layer 3. The visualization does not display any meaningful clustering prior to the tagging mechanism.
(b) Attention probabilities at layer 3, head 30.
(c) Activation outliers in attention output at layer 3, head 30.
(d) Value embedding at layer 3, head 30.
(e) PCA on residual stream at layer 18.
Figure 7: Visualization of the ‘catch, tag, release’ mechanism on a sample prompt.
(a) PCA on the input embeddings to attention layer 3.
(b) Attention weights of layer 3, head 6.
(c) Output of attention layer 3, head 6.
(d) PCA on the output of attention layer 3, head 6.
(e) PCA on the residual stream at layer 3 after the attention layer.
(f) Attention sink at layer 38, head 33,
highlighting “Let’s think step by step”.
(g) PCA on the residual stream at layer 38, preserving earlier tagging effects.
Figure 8: Visualization of the ‘catch, tag, release’ mechanism in Zero-Shot Chain of Thought (CoT) prompt. Later layers reveal an attention sink forming on “Let’s think step by step”, with the output embedding reflecting earlier tagged value embeddings.
(a) PCA on the input embeddings to attention layer 5.
(b) Attention weights of layer 5, head 16.
(c) Output of attention layer 5, head 16.
(d) PCA on the output of attention layer 5, head 16.
(e) PCA on the residual stream at layer 21.
Figure 9: Visualization of the ‘catch, tag, release’ mechanism on a Chain of Thought (CoT) prompt.
(a) PCA on the input embeddings to attention layer 5.
(b) Attention weights at layer 5, head 14.
(c) Output of attention layer 5, head 14.
(d) PCA on the output of attention layer 5, head 14.
(e) PCA on the residual stream at layer 18.
(f) Attention sink at layer 39, head 24,
highlighting “Let’s think step by step”.
(g) PCA on residual stream at layer 39, preserving earlier tagging effects.
Figure 10: Visualization of the ‘catch, tag, release’ mechanism on a Zero-Shot Chain of Thought (CoT) prompt. Later layers reveal an attention sink forming on ”Let’s think step by step”, with the output embedding retaining features from tagged value embeddings.
(a) PCA on the input embeddings to attention layer 3.
(b) Attention weights of layer 3, head 30.
(c) Output of attention layer 3, head 30.
(d) PCA on the output of attention layer 3, head 30.
(e) PCA on the residual stream at layer 16.
Figure 11: Visualization of the ‘catch, tag, release’ mechanism on a Chain of Thought (CoT) prompt.
Appendix C Experiment Details
We use the HuggingFace Transformers library to load the models utilized in our experiments [Wolf et al., 2019].
C.1 Pruning Hyperparameters
For all pruning experiments, we prune uniformly across all linear layers in the model, excluding the embedding and head layers, remaining consistent with [Frantar and Alistarh, 2023, Sun et al., 2024b, Zhang and Papyan, 2024]. As calibration data, we use 128 sequences of length 2048 from the first shard of the C4 dataset [Raffel et al., 2019]. The algorithm-specific hyperparameters are:
  • SparseGPT
    • Hessian Dampening: $0.1$
    • Block Size: $128$
  • OATS
    • Iterations: $80$
    • Rank Ratio: $0.25$
The prompt used to generate Figure 5 is:
[…]
which we sourced from Yu et al. [2024].

### v2

@@ D.2 Interaction Between Tags and Embeddings
E.1 Proof of Theorem, Part 1: ‘Catch, Tag, Release’
@@
Thus, [SEP]acts as an attention sink, where all tokens $x_{i}$ for $i>t$, as well as the [SEP] token itself, attend exclusively to the [SEP]token. As a result, the attention of all tokens $x_{i}$ for $i>t$ and the [SEP] token have been caught.
@@
$=$ ${\bm{A}}^{1}_{i,t}\begin{bmatrix}0&s_{\texttt{tag}}\end{bmatrix}+\sum_{j=1,j\neq t}^{i}{\bm{A}}^{1}_{i,j}{\bm{e}}_{j}^{\top}{\bm{W}}^{1}_{V}\stackrel{{\scriptstyle s_{\texttt{tag}}\to\infty}}{{\xrightarrow{\hskip 28.45274pt}}}\begin{bmatrix}0&\infty\end{bmatrix}$.
@@
E.2 Proof of Theorem, Part 2: Leveraging the Tags
@@
$\exp({\bm{h}}_{T}^{\top}{\bm{W}}^{2}_{Q}{\bm{W}}^{2}_{K}({\bm{h}}_{j}-{\bm{h}}_{T}))\stackrel{{\scriptstyle s_{\texttt{tag}}\to\infty}}{{\xrightarrow{\hskip 28.45274pt}}}0.$
@@
$\exp({\bm{h}}_{T}^{\top}{\bm{W}}^{2}_{Q}{\bm{W}}^{2}_{K}({\bm{h}}_{t}-{\bm{h}}_{T}))\stackrel{{\scriptstyle s_{\texttt{tag}}\to\infty}}{{\xrightarrow{\hskip 28.45274pt}}}0.$
@@
$\exp({\bm{h}}_{T}^{\top}{\bm{W}}^{2}_{Q}{\bm{W}}^{2}_{K}({\bm{h}}_{j}-{\bm{h}}_{T}))\stackrel{{\scriptstyle s_{\texttt{tag}}\to\infty}}{{\xrightarrow{\hskip 28.45274pt}}}1.$
Combining all three cases together with Equation (8) leads to:
@@
The above shows how the tags have been leveraged to arrive at a uniform distribution over the desired tokens. Then using this, Equation (7), and that ${\bm{W}}^{2}_{V}=\begin{bmatrix}1\\ 0\end{bmatrix}$, we can conclude that:
@@
E.3 Proof of Theorem, Part 3: Proving the Three Cases
@@
where the order of magnitude of the last two terms is given by Equation (5).
@@
which we show in Appendix E.4.
@@
where $\lim_{s_{\texttt{tag}}\to\infty}({\bm{A}}^{1}_{j,t}-{\bm{A}}^{1}_{T,T})s_{\texttt{tag}}=0$ is proved in Appendix E.4 below.
E.4 Limit Proof
@@
$\lim_{s_{\texttt{tag}}\to\infty}s_{\texttt{tag}}^{2}({\bm{A}}^{1}_{t,t}-{\bm{A}}^{1}_{T,t})$
$=\lim_{s_{\texttt{tag}}\to\infty}\frac{\sum_{k=1,k\neq t}^{T}s_{\texttt{tag}}^{2}\exp(x_{T}x_{k}+1-s_{\texttt{tag}})-(t-1)s_{\texttt{tag}}^{2}\exp(s_{\texttt{tag}}-s_{\texttt{tag}}^{2})}{\left(1+(t-1)\exp(s_{\texttt{tag}}-s_{\texttt{tag}}^{2})\right)\cdot\left(1+\sum_{k=1,k\neq t}^{T}\exp(x_{T}x_{k}+1-s_{\texttt{tag}})\right)}$
$=\frac{0}{1}=0$
@@
$\lim_{s_{\texttt{tag}}\to\infty}({\bm{A}}^{1}_{j,t}-{\bm{A}}^{1}_{T,T})s_{\texttt{tag}}$
$=\frac{\sum_{k=1,k\neq t}^{T}s_{\texttt{tag}}\exp(x_{T}x_{k}+1-s_{\texttt{tag}})-\sum_{k=1,k\neq t}^{j}s_{\texttt{tag}}\exp(x_{j}x_{k}+1-s_{\texttt{tag}})}{\left(1+\sum_{k=1,k\neq t}^{j}\exp(x_{j}x_{k}+1-s_{\texttt{tag}})\right)\cdot\left(1+\sum_{k=1,k\neq t}^{T}\exp(x_{T}x_{k}+1-s_{\texttt{tag}})\right)}$
@@
$\lim_{s_{\texttt{tag}}\to\infty}({\bm{A}}^{1}_{j,t}-{\bm{A}}^{1}_{T,T})s_{\texttt{tag}}^{2}$
$=\frac{\sum_{k=1,k\neq t}^{T}s_{\texttt{tag}}^{2}\exp(x_{T}x_{k}+1-s_{\texttt{tag}})-\sum_{k=1,k\neq t}^{j}s_{\texttt{tag}}^{2}\exp(x_{j}x_{k}+1-s_{\texttt{tag}})}{\left(1+\sum_{k=1,k\neq t}^{j}\exp(x_{j}x_{k}+1-s_{\texttt{tag}})\right)\cdot\left(1+\sum_{k=1,k\neq t}^{T}\exp(x_{T}x_{k}+1-s_{\texttt{tag}})\right)}$
@@
Appendix F Sensitivity Analysis of the Sink Threshold
We performed a comprehensive sensitivity analysis by varying the threshold across twelve values from $0.03$ to $0.4$ and evaluating its effect on two key metrics:
  1. 1.
    the average number of identified sinks, and
  2. 2.
    the average variance explained by the tags.
The results are shown in Figure 21 above.
1-10112233445566778899101000.10.10.20.20.30.30.40.40.50.5Avg. Number of SinksAvg. Explained VarianceDeepSeek LLaMA 8bDeepSeek Qwen 14BDeepSeek Qwen 7BLLaMA 3.1 8BLLaMA 3 8BPhi-3 MediumPhi-3 MiniPhi-4Qwen 2.5 14BQwen 2.5 3BQwen 2.5 7BQwen 3 14BQwen 3 8B
Figure 21: Average explained variance plotted against the average number of sinks for different values of $\varepsilon$. Each marker indicates a specific $\varepsilon$ in the set $\{0.03,0.04,0.05,0.06,0.07,0.08,0.09,0.1,0.15,0.2,0.3,0.4\}$.
Across all models tested, the resulting curves are typically logarithmic-like: initially, decreasing the threshold adds new sinks and increases explained variance, but this gain saturates beyond a certain point. This suggests that smaller thresholds tend to capture less meaningful or redundant tokens. The table below provides representative values from this analysis to support this observation.
Appendix G Optimization Details for Section 7.3
Figure 22: The predicted averages of the trained model versus the actual average on the evaluation set.
We generate a total of 16,384 sequences of length 16 where half are allocated for training while the other half is allocated for evaluation. For each sequence, a random index is generated to place the [SEP]. We minimize the mean squared error loss between the predicted and actual average and employ a cosine annealing learning rate scheduler. We employ an initialization of $s_{\texttt{tag}}=10$ to aid with convergence. Depicted in Figure 22 are the predicted averages of the trained model versus the actual average on the evaluation set.
Appendix H Qwen 3 Results
We present quantitative evidence that the Qwen 3 models [Yang et al., 2025], which incorporate QK normalization, exhibit the ‘catch, tag, release’ mechanism.
H.1 Sink Count and Variance Explained by Sinks
(a) Attention Sink Counts for Qwen 3. Counts are computed for each head and their distribution is visualized for each model using violin plots.
(b) Variance Explained by Tags for Qwen 3. The metric is computed for each head and summary statistics are shown for each model using box plots.
Figure 23: Quantitative Measurements for Qwen 3. Analogous measurements presented in Section 2 for models in the Qwen 3 family. Notably, these models exhibit the ‘catch, tag, release’ mechanism while having QK normalization.
H.2 Probe Performance
Probe Qwen 3
8B 14B
$\theta_{\text{tag}}$ $\mathbf{100\%}$ $\mathbf{87.0\%}$
$\theta_{\text{no tag}}$ $50.5\%$ $64.5\%$
$\theta_{\text{activation}}$ $56.0\%$ $82.0\%$
Table 5: Classification Accuracy of Probes for Qwen 3. The probe $\theta_{\text{tag}}$ is computed from the tag component of the activation, $\theta_{\text{no tag}}$ from the non-tag component, and $\theta_{\text{activation}}$ from the full activation.
Appendix I Additional Discussion
I.1 Practical Applications to LLMs
The presence and utility of the ‘catch, tag, release’ mechanism have important implications for practical use of large language models. Section 4 shows the potential for tags to be used to steer activations [Subramani et al., 2022, Turner et al., 2025] which can be applied for model alignment and safety [Wang and Shu, 2024]. In prior work Yona et al. [2025], attention sinks were linked to repeated token phenomena and jailbreaking, suggesting the potential for the mechanism to be used for both exploit detection and mitigation. In model compression, preserving this mechanism may be critical, as we hypothesize it can be captured within a low-rank subspace of the model’s parameters, which must be retained during pruning or quantization to maintain performance [Zhang and Papyan, 2025b, Makni et al., 2025]. Furthermore, recent work, Wang et al. [2025], has identified low-entropy tokens as central to LLM reasoning abilities, likely overlapping with the tag tokens we highlight, indicating this mechanism may play a foundational role in LLM reasoning.
I.2 Role of the Fully Connected Layers
Empirically, prior work has shown that groups of token representations tend to collapse to low-dimensional subspaces, or even single points, across layers [Geshkovski et al., 2025]. This raises the question: which tokens collapse together, and what determines the target of this collapse? We hypothesize that the ‘catch, tag, release’ mechanism plays a central role in both the grouping and the collapse target. Specifically, tokens attending to the same attention sink receive a common tag, which defines the direction and destination of their collapse. We further propose that the MLP layers, via a low-rank subspace, progressively drive this collapse across layers by reinforcing the shared tag structure.
Appendix J Mass-Mean Probe Details
The ([CITY], [COUNTRY]) pairs are sourced from GeoNames [GeoNames, 2025].
J.1 Layer and Head Information
For Table 1, the layer and attention head for each model that we extracted the probes from are depicted in Table 6 below:
Qwen 2.5 Llama-3 8B Llama-3.1 8B
3B 7B 14B
Layer $24$ $16$ $40$ $18$ $18$
Head 44 $28$ $40$ 2121 2121
Table 6: Layer/Attention Head information for results presented in Table 1.
J.2 Taxonomy of Tag Tokens
For each tag-only probe presented in Table 1, we present the histogram of the tag tokens that emerged during their generation in Figure 24 below.
Qwen 2.5 3B
Qwen 2.5 7B
Qwen 2.5 14B
LLaMA-3 8B
LLaMA-3.1 8B
Figure 24: A taxonomy of the tokens that were identified as attention sinks during the generation of the $\theta_{\text{tag}}$ and $\theta_{\text{no tag}}$ probes. [PREAMBLE] is being used to represent all tokens that correspond to the following portion of the prompt: The city of Tokyo is in Japan. This statement is TRUE. The city of Hanoi is in Poland. This statement is: FALSE.
J.3 Definition of $\theta_{\text{activation}}$
Following Marks and Tegmark [2024], the activation-based probes, $\theta_{\text{activation}}$, are computed from the full activations:
${\bm{z}}_{t}=\sum_{k=1}^{T}{\bm{A}}_{t,k}{\bm{V}}_{k,:}.$
From these representations, we obtain the class means:
$\mu^{+}=\frac{1}{N_{+}}\sum_{i\in\text{True}}{\bm{z}}_{t}^{(i)},\quad\mu^{-}=\frac{1}{N_{-}}\sum_{i\in\text{False}}{\bm{z}}_{t}^{(i)}.$
The within-class covariance matrix, $\Sigma$, is then estimated, and the resulting mass-mean probes are defined as:
$\theta_{\text{activation}}({\bm{z}})=\sigma\!\left({\bm{z}}^{\top}\Sigma^{-1}(\mu^{+}-\mu^{-})\right).$
J.4 Sentiment Analysis
it’s a charming and often affecting journey. This statement is: POSITIVE. unflinchingly bleak and desperate. This statement is: NEGATIVE. [SENTENCE]. This statement is:
Figure 25: Sentiment prompt example.
To explore how the tagging mechanism can potentially capture a broader range of semantic and syntactic information beyond just binary True/False classifications, we extend our experiments to assess whether tags can encode sentiment. Specifically, we construct prompts in the format shown in Figure 25 on the right.
[SENTENCE] is replaced with various examples from the SST-2 dataset [Socher et al., 2013]. Using the same experimental setup described in Section 4, we evaluated whether tags could capture sentiment polarity by decomposing the activations of the final period token (indicated by the arrow).
The results, presented in Table 7 below, show that tags indeed encode positive and negative sentiment information effectively, supporting the broader applicability of the method to a variety of different semantic tasks.
Probe Qwen 2.5 LLaMA-3 LLaMA-3.1
3B 7B 14B 8B 8B
$\theta_{\text{tag}}$ $\mathbf{88.5\%}$ $\mathbf{90.0\%}$ $\mathbf{94.0\%}$ $\mathbf{87.5\%}$ $\mathbf{86.5\%}$
$\theta_{\text{no tag}}$ $84.0\%$ $53.0\%$ $53.5\%$ $81.5\%$ $64.0\%$
$\theta_{\text{activation}}$ $86.5\%$ $80.0\%$ $90.5\%$ $88.0\%$ $83.0\%$
Table 7: Classification Accuracy of Probes. Comparison of the three probe sets on positive/negative sentiment analysis.
Appendix K Limitations
K.1 Theoretical Model
While our theoretical result provides a constructive proof that the ‘catch, tag, release’ mechanism can emerge in transformer architectures, it comes with limitations detailed below.
Theorem 7.1 does not prove that the ‘catch, tag release’ mechanism is necessary. There may exist alternative solutions that perform the task without attention sinks or outlier features. Consequently, the theoretical result should be viewed as illustrative. Although we provide empirical support suggesting that similar dynamics arise in trained models, we do not prove convergence to our specific construction under training.
The theoretical model operates under highly constrained conditions: a two-layer transformer solving a sequence averaging problem. While this setting is valuable for analytical tractability, it is removed from the complexity of real-world language modeling tasks.
Furthermore, depite the theoretical model being simple, the optimization process does not always reach a low-loss solution. We found that convergence depends on how the $s_{\texttt{tag}}$ parameter is initialized, with larger initial values generally leading to more consistent success. Table 8 below shows how often the model achieved a high fit ($R^{2}>0.95$) across 10 runs for different starting values of $s_{\texttt{tag}}$:
$s_{\text{tag}}$ initialization Success Rate
10 4/10
6 3/10
1 0/10
Table 8: Success rate, measured by $R^{2}>0.95$, across different $s_{\text{tag}}$ initializations.
K.2 Connection with Reasoning
While our findings in Section 5 strongly suggest a structural role for sinks in reasoning, they do not yet establish a causal link between the two. This remains an important direction for future work.
K.3 Threshold-Based Metric
The threshold-based metric used to identify attention sinks does not guarantee full coverage of all such tokens. Our choice of a $\varepsilon=0.2$ threshold, while consistent with prior studies, is not necessarily optimal. As discussed in Gu et al. [2024], there is currently no principled method for determining this threshold. A sensitivity analysis for this threshold is provided in Appendix F.
Appendix L Compute Resources
All experiments involving LLMs were executed utilizing a single NVIDIA A40 with 48GB of GPU memory. This includes the experiments used to generate Figures 2, 4, 6, 7, Table 1, and visualizations provided in Appendix A. Each of the experiments in Sections 4, 6, 7 took approximately $40$ minutes per model while the visualizations in Section 2 took roughly $60$ minutes per model totalling for roughly 40 hours. Roughly 30 additional hours were utilized for preliminary experiments that did not reach the final paper.
Appendix M Models and Implementation
A comprehensive list of all the models used in our empirical study is found below:
  • Phi-3 Mini and Phi-3 Medium [Abdin et al., 2024a]
  • Phi-4 [Abdin et al., 2024b]
  • Qwen 2.5 3B, 7B, 14B, 32B [Yang et al., 2024]
  • LLaMA-3 8B, LLaMA 3.1 8B [Grattafiori et al., 2024]
  • Mistral 7B v0.3 [Jiang et al., 2023]
  • Deepseek-R1 LLaMA 8B, Deepseek-R1 Qwen 14B, Deepseek Math 7B [DeepSeek-AI, 2025]
  • Gemma 2 9B [Gemma Team et al., 2024]
  • Gemma 3 4B, 12B [Gemma Team et al., 2025]
  • Qwen 3 4B, 8B, 14B [Yang et al., 2025]
We utilize the Transformers library by Huggingface [Wolf et al., 2020] to run all LLM experiments.